Find an equation of the line described below. Write the equation in slope-intercept form (solved for y), when possible.
Through (13,9) and (9,13)
What is the equation of the line?
(simplify your answer.)

Answer by mananth(16949)

equation of line between two points
m= 13-9/9-13
=-1
y-y1=m(x-x1)
y-9=-1(x-13)
y-9=-x+14
y=-x+23
...
m.ananth@hotmail.ca